When the rate of disappearance of O2 is 0.13 Ms−1, the rate of appearance of CO2 is ________ Ms−1.
Answer #1

THE REACTION IS NOT PROVIDED.

SO, I am using this reaction,

C2H4 (g)+3O2(g)→2CO2(g)+2H2O(g) [BALANCED]

Rate = change in Concentration/change in time

For a General reation,

aA + bB \rightarrow cC + dD

Rate = -(1/a)*(\Delta[A]/\Deltat) = -(1/b)*(\Delta[B]/\Deltat ) = (1/c)*(\Delta[C]/\Deltat) = (1/d)*(\Delta[D]/\Deltat)

Negative sign is just to show that they are disappearing

So, For our reaction,

C2H4 (g)+3O2(g)→2CO2(g)+2H2O(g)

Rate = -(\Delta[C2H4]/\Deltat) = -(1/3)*(\Delta[O2]/\Deltat ) = (1/2)*(\Delta[CO2]/\Deltat) = (1/2)*(\Delta[H2O]/\Deltat) ---------------A

ANSWER -

Given

rate of disappearance of O2 = 0.13 M/s

Rate of appearance of CO2 = ?

From A,

-(1/3)*(\Delta[O2]/\Deltat ) = (1/2)*(\Delta[CO2]/\Deltat)

rate of appearance of CO2 (\Delta[CO2]/\Deltat)= (2/3) * rate of disappearance of O2 (\Delta[O2]/\Deltat)

Put the values,

rate of appearance of CO2 = (2/3) * rate of disappearance of O2

rate of appearance of CO2 = (2/3) * 0.13 M/s

rate of appearance of CO2 = 0.0867 M/s [ANSWER]